Before After (2.90 ku louity of ,90 mV 00 k the two possible answers two this problem, choose the solution which has ml going as fast as possible) What is the angle phi, between mass twos valority and the initial clarity nr mass nna? ive ynur answer as a positive number in degrees.) Submit Men two after th 31.0 de peed of Hole, the
Answer #1

m1 = 2.80kg ļ Vo = 5.90m 15 m2 = 2kg (at rest initially) v= 95mm > consonation of KE C collision is elastic ) men 1+ 1 = + ,fueritte Larango €9 @ >> Ve - mive 1958 m2 V (60 - 4 Noces squarury 4 odsung eq?? ☺ f@ => [mive Sine ) 2 + cm, Vo - , Leo) 2Vi= 79.2982 6 1 45.78 66 2 X 13.43967 Vis 40 65 36 & 1.2467 ::V, (max) = 406536 mis : from o = m, no? = mui2 + M₂ V ₂ 2 =) mi